If I add water to 100 ml of a 0.15 M NaOH solution until the final volume is 150 ml, what is the molarity of the diluted solution be?
1
Expert's answer
2021-10-01T11:13:01-0400

Molar mass of NaOH = 39.997

0.15×39.997 = 5.996

100/5.996= 16.68g

16.68/39.997 = 0.417mol

Molarity = 0.417/0.15

= 2.78mol/L